OpenStudy (nopeforce):

What is the slope of the line that passes through the points (–2, 1) and (1, 6)? a.-5 b.-7 c.3/5 d.5/3 Im positive its D

rebeccaxhawaii (rebeccaxhawaii):

graph it

alones (alones):

https://www.desmos.com/

pooja195 (pooja195):

\[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \] \[ \huge~\rm~\bf\color{blue}{ \frac{6-1=5}{1--2}=slope} \] \[ \huge~\rm~\bf\color{blue}{ \frac{6-1=5}{1+2=3}=slope} \] so correct

alones (alones):

Grpah it there click it :P

pooja195 (pooja195):

Why not just help instead?

pooja195 (pooja195):

Instead of taking the lazy way out.

alones (alones):

Pooja were ... :3

rebeccaxhawaii (rebeccaxhawaii):

tru ^^

OpenStudy (nopeforce):

Yeah thx, I already graphed it

alones (alones):

mhm so pooja don't count on us